OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = 9x + 3 and g(x) = Quantity x minus three divided by nine

OpenStudy (rogue):

Simple do as the question says :) Show that f(g(x)) = x and g(f(x)) = x. In order to show that g(f(x)) = x, plug in f(x) into the x of g(x). So for example, f(x) = 9x + 3 and g(x) = (x - 3)/9 f(x) ----> g(x) g(f(x)) = (f(x) - 3)/9 = ((9x + 3) - 3) / 9 = (9x)/9 = x Now you have to do the second half by showing that f(g(x)) = x. You can follow the same steps for g(f(x)).

OpenStudy (anonymous):

ohhh I see now. Thank you so much for your time ^_^